### v2

@@ Abstract
Electronic Health Records (EHRs) offer rich potential for clinical prediction, yet their inherent complexity and heterogeneity pose significant challenges for traditional machine learning approaches. Domain-specific EHR foundation models trained on large collections of unlabeled EHR data have demonstrated promising improvements in predictive accuracy and generalization; however, their training is constrained by limited access to diverse, high-quality datasets and inconsistencies in coding standards and healthcare practices. In this study, we explore the possibility of using general-purpose Large Language Models (LLMs) based embedding methods as EHR encoders. By serializing patient records into structured Markdown text, transforming codes into human-readable descriptors, we leverage the extensive generalization capabilities of LLMs pretrained on vast public corpora, thereby bypassing the need for proprietary medical datasets. We systematically evaluate two state-of-the-art LLM-embedding models, GTE-Qwen2-7B-Instruct and LLM2Vec-Llama3.1-8B-Instruct, across 15 diverse clinical prediction tasks from the EHRSHOT benchmark, comparing their performance to an EHR-specific foundation model, CLIMBR-T-Base, and traditional machine learning baselines. Our results demonstrate that LLM-based embeddings frequently match or exceed the performance of specialized models, even in few-shot settings, and that their effectiveness scales with the size of the underlying LLM and the available context window. Overall, our findings demonstrate that repurposing LLMs for EHR encoding offers a scalable and effective approach for clinical prediction, capable of overcoming the limitations of traditional EHR modeling and facilitating more interoperable and generalizable healthcare applications.
1 Introduction
@@
EHRs have become a widespread technology in modern healthcare, providing a comprehensive, longitudinal view of a patient’s health status [1]. Machine learning methods can leverage this rich data to perform risk stratification and support clinical decision making [2, 3, 4]. In recent years, researchers have explored a variety of prediction tasks based on EHRs, including hospital readmission [5, 6], length of hospital stay [6], sepsis onset detection [7, 8], mortality prediction [6, 9], discharge diagnoses [6], and heart failure outcomes [10]. The overarching objective is to harness existing EHR data through machine learning to enhance clinical outcomes and reduce healthcare costs.
@@
However, machine learning on EHR data poses significant challenges due to its inherent complexity. EHR data is characterized by variable-length sequences of patient visits, irregular sampling intervals, missing entries, heterogeneous and noisy information, and a wide range of hierarchical medical concepts [11]. As a result, deep learning models often achieve only modest improvements over traditional methods such as logistic regression or tree-based methods [6, 12, 13]. To mitigate these issues, recent approaches have employed large-scale foundation models that are pre-trained on unlabeled EHR data using unsupervised learning [14]. Many of these models adopt strategies from natural language processing, such as masked-word prediction as in BERT [15] or autoregressive next-word prediction as in GPT [16]. Treating EHR data as sequences of medical codes, enables analogous methods such as masked code prediction [17, 18, 12, 19] or next code prediction [13]. However, these techniques also face significant limitations: coding standards and healthcare practices differ strongly across sites, and interoperable EHR foundation models would likely need to be trained on a wide variety of EHR datasets, which is difficult to achieve due to the sensitivity of healthcare data. Therefore, the development of EHR-specific foundation models remains constrained by the limited size and restricted availability of EHR data.
@@
In contrast, LLMs benefit from training on vast general-purpose text corpora and a broad range of natural-language tasks [20]. This extensive pre-training enables their language comprehension and allows them to capture domain-agnostic patterns that can be adapted for healthcare applications. Consequently, LLMs have demonstrated strong performance in extracting medical concepts [21], summarizing medical texts [22], and predicting medical outcomes [23] even in low-resource settings. However, most modern LLMs, such as GPT [24] or Llama [25], utilize a decoder-only transformer architecture, which complicates the generation of robust text representations. To overcome this limitation, recent work has introduced methods to convert decoder-only LLMs into effective embedding models for downstream prediction tasks [26, 27, 28, 29]. Additionally, these state-of-the-art models offer an increased context window, making them well-suited for handling long inputs such as serialized EHR data.
@@
In this study, we systematically evaluate whether general-purpose LLM-embedding models can effectively encode EHR records for 15 distinct clinical prediction tasks [30] (see Fig. 1). To this end, we first transform the EHR data into a concise text representation capturing the most relevant patient information at prediction time. We then use two LLM-embedding models, GTE-Qwen2-7B-Instruct [31, 29] and LLM2Vec-Llama3.1-8B-Instruct [26, 32], to generate EHR representations. The resulting embeddings serve as inputs to a logistic regression classifier, which is trained separately for each prediction task. We focus on evaluating the performance in few-shot settings to evaluate the generalization ability of this approach and conduct extensive ablation studies to identify the specific components that drive the LLM ’s effectiveness.
@@ 2 Results
2.1 EHR Database and Prediction Tasks
We used the EHRSHOT database containing adult patients from the Stanford Health Care and Lucile Packard Children’s Hospital from 1990 to February 8th, 2023 [30]. This dataset includes comprehensive EHR timelines for 6.739 patients, covering 921.499 visits, and 41.661.637 clinical events (Table 1). The database is part of a rigorous EHR benchmark containing 15 clinical prediction tasks grouped into four task groups. It also provides canonical dataset splits and provides publicly available code. Table 2 presents a detailed breakdown of the task groups, individual tasks, and the corresponding number of labels used in our experiments. Notably, a single patient may contribute multiple labels for a given prediction task, as relevant clinical events can recur over time.
| Attribute | Train | Validation | Test | Total |
| Num Events | 15,511,472 (37.2) | 13,005,205 (31.2) | 13,144,960 (31.6) | 41,661,637 (100.0) |
| Num Visits | 339,504 (36.8) | 300,325 (32.6) | 281,670 (30.6) | 921,499 (100.0) |
| Num Patients | 2,295 (34.1) | 2,232 (33.1) | 2,212 (32.8) | 6,739 (100.0) |
| Num Female | 1,173 (34.1) | 1,142 (33.2) | 1,126 (32.7) | 3,441 (100.0) |
| Num Male | 1,122 (34.0) | 1,090 (33.1) | 1,086 (32.9) | 3,298 (100.0) |
| Age, mean SD | 59.6 17.8 | 58.8 18.0 | 59.5 18.0 | 59.3 17.9 |
| American Indian | 14 (56.0) | 7 (28.0) | 4 (16.0) | 25 (100.0) |
| Asian | 356 (34.1) | 347 (33.3) | 340 (32.6) | 1,043 (100.0) |
| Black | 98 (32.9) | 105 (35.2) | 95 (31.9) | 298 (100.0) |
| Pacific Islander | 23 (31.1) | 21 (28.4) | 30 (40.5) | 74 (100.0) |
| Unknown | 518 (33.1) | 530 (33.9) | 515 (32.9) | 1,563 (100.0) |
| White | 1,286 (34.4) | 1,222 (32.7) | 1,228 (32.9) | 3,736 (100.0) |
| Hispanic | 374 (36.0) | 342 (32.9) | 322 (31.0) | 1,038 (100.0) |
| Non-Hispanic | 1,921 (33.7) | 1,890 (33.2) | 1,890 (33.2) | 5,701 (100.0) |
@@ 2.1 EHR Database and Prediction Tasks
2.2 EHR Text Serialization
To utilize an LLM-embedding model for EHR representation, the patient records were encoded into a structured text based on the widely used Markdown format. Due to runtime constraints, the maximum serialization length was limited to 4.096 tokens (approximately 16.000 characters). This constraint informed the serialization strategy, prioritizing the inclusion of recent data to ensure critical medical information was preserved. The serialized record is divided into clearly labeled sections (see example in Fig. 2). All dates were normalized relative to a reference prediction date of January 1st, 2024, explicitly stated at the beginning. This is followed by the patient’s basic demographic information. Approximately 65% of the recorded values were time-series data of Logical Observation Identifiers Names and Codes (LOINC) concepts. To reduce the volume of this data, the serialization focuses on 24 frequently recorded concepts, grouped into three primary categories: Body Metrics, Vital Signs, and Lab Results. For each selected concept, the last three recorded values were included, along with the corresponding units and a classification as low, normal, or high, where applicable. The serialization then summarizes all patient visits, including visit type, time and duration, followed by any event not associated with specific visits. Detailed visit records are then provided in descending chronological order, beginning with the most recent visit. Each visit entry is further categorized into conditions, medications, and procedures for improved clarity and utility. As a result, the text serialization ensures a structured and concise representation of EHR data, facilitating efficient processing by LLM-based models.
2.3 LLM-Embedding Models and Baselines
We focus on two LLM-embedding models as part of our experiments to explore their effectiveness in representing EHR data. The first model, GTE-Qwen2-7B-instruct (GTE-Qwen2-7B), is based on the Qwen2-7B LLM [31] and incorporates bidirectional attention and contrastive learning to enhance embedding tasks [29]. The second model, LLM2Vec-Llama-3.1-8B-Instruct (LLM2Vec-Llama-3.1-8B), is built on the Llama-3.1-8B Instruct architecture [32] and employs similar optimization techniques to improve embeddings [26]. Both models were trained using instructions for the embeddings task. Hence, we added a simple prompt for each task, e.g., “Given a patient’s electronic healthcare record (EHR) in Markdown format, retrieve relevant passages that answer the query: has the patient anemia?” (see Table 8). Our primary goal was to assess how these LLM-based models, trained on publicly available text data, perform in representing EHRs compared to an EHR-specific foundation model. For this, we included CLMBR-T-Base, a 141-million-parameter autoregressive foundation model trained on 2.57 million de-identified EHRs from Stanford Medicine [13, 30]. For each model, we used a logistic regression classification head trained on the training split, with hyperparameters tuned on the validation split. Additionally, we included a counts-based baseline using a Gradient Boosted Machine (GBM) model, which has demonstrated superior performance over logistic regression for EHR tasks [30].
@@ 2.3 LLM-Embedding Models and Baselines
2.4 Performance Results on 15 Prediction Tasks
@@
The performance results measured by the Area Under the Receiver Operating Characteristic Curve (AUROC) for all training and validation examples are presented in Table 3. The GTE-Qwen2-7B model demonstrated superior performance over CLIMBR-T-Base in three of the four task categories: operation outcomes, lab test result prediction, and assignment of new diagnoses. However, in predicting chest X-ray findings, the EHR foundation model (CLIMBR-T-Base) outperformed GTE-Qwen2-7B. The LLM2Vec-Llama-3.1-8B embeddings showed slightly lower performance compared to GTE-Qwen2-7B, surpassing CLIMBR-T-Base only in the assignment of new diagnoses task. Both LLM-based embedding models clearly outperformed the counts-based baseline with a gradient-boosted machine head, which is traditionally regarded as a strong baseline for EHR tasks [12, 13]. The only exception was in the assignment of new diagnoses, where the counts baseline marginally outperformed GTE-Qwen2-7B. We also combined the CLMBR-T-Base embedding with the LLM representations by simple concatenation to test whether they encode orthogonal information. The combined embeddings led to a considerable performance boost for both models with an average AUROC performance of 0.801 (0.777 - 0.826) instead of 0.774 (0.749 - 0.798) for GTE-Qwen2-7B and 0.779 (0.753 - 0.805) instead of 0.742 (0.714 - 0.769) for LLM2Vec-Llama-3.1-8B. The smaller GTE-Qwen2-1.5B model only performed slightly worse than GTE-Qwen2-7B with an average performance of 0.757 (0.732 - 0.783) compared to 0.774 (0.749 - 0.798) making it a performant model choice for EHR embeddings. LLM2Vec-Llama-2-1.3B on the other hand performed significantly worse than the larger model which could be due to the older Llama 2 architecture. Fig. 3 illustrates the relationship between model size and average performance across all tasks, suggesting the trend that larger models tend to perform better. Among the evaluated models, CLIMBR-T-Base proved to be the most efficient in terms of performance relative to model size. It is noteworthy, however, that the LLM-based embedding models were not specifically optimized for this EHR data modality, which may partially explain their performance gap.
2.5 Performance Results in Few-Shot Setting
@@
To evaluate model performance with limited training data, we conducted experiments in a few-shot setting using small numbers of training examples. The experiments followed the EHRSHOT task definitions [30]. Fig. 4 illustrates the aggregated AUROC across all subtasks within the four task categories for varying numbers of training examples (see Fig. 6 for AUPRC results). Both LLM-embedding models demonstrated strong results in the few-shot setting, indicating that their pretraining on general text can be successfully transferred to serialized EHR data. Notably, GTE-Qwen2-7B consistently outperformed CLIMBR-T-Base across all training example sizes for predicting lab test results and assigning new diagnoses. For operational outcomes, a minimum of 32 training examples was required for GTE-Qwen2-7B to surpass the performance of the EHR-specific foundation model. In contrast, LLM2Vec-Llama-3.1-8B showed weaker performance in the few-shot setting. It only outperformed CLIMBR-T-Base in a limited number of scenarios, primarily for the assignment of new diagnoses. Both LLM-embedding models consistently achieved improvements over the counts-based baseline across most configurations, underscoring the impact of their pretraining on general text. Consistent with previous findings [30], the largest performance gains over the counts-based baseline were observed at intermediate training example sizes. As the number of training examples increased, the advantage of pretrained LLM-based models decreased, highlighting the diminishing returns of foundation models as more labeled data becomes available.
2.6 Effect of Different Contexts Sizes
We evaluated the impact of varying context sizes on the performance of the LLM-embedding models by testing GTE-Qwen2-7B and LLM2Vec-Llama-3.1-8B with input lengths of 512, 1.024, 2.048, 4.096 (the default), and 8.192 tokens. The objective was to determine whether longer input sequences enhance performance and to assess if the models effectively leverage additional historical data. As summarized in Table 4 and Fig. 5, the two models exhibited distinct behaviors.
@@ 2.6 Effect of Different Contexts Sizes
| 8.192 context size | |||||
| 2.048 context size | |||||
| 1.024 context size | |||||
| 512 context size | |||||
| LLM2Vec-Llama-3.1-8B | |||||
| 8.192 context size | |||||
| 2.048 context size | |||||
| 1.024 context size | |||||
| 512 context size |
For GTE-Qwen2-7B, the best overall performance was achieved with a context of 4.096 tokens, with only a slight decline observed when using 2.048 or 1.024 tokens. This suggests that the most relevant information is contained within the first 1.024 tokens and that the model can accurately extract these details even when presented within a longer context. In contrast, LLM2Vec-Llama-3.1-8B showed a marked improvement with a context size of 2.048 tokens and a moderate gain at 1.024 tokens, indicating that the model struggles to effectively handle longer sequences up to 4.096 tokens. Both models experienced a significant drop in performance when limited to only 512 tokens, highlighting that such a short context omits crucial information; a noteworthy point given that many existing language models are constrained to 512 input tokens [33]. Furthermore, extending the context to 8.192 tokens resulted in decreased performance for both models, implying limitations in extracting relevant information from very long inputs. This decline may be partially attributed to the mean pooling of the last hidden layers used for generating embeddings, which can dilute the impact of important token-level information; an alternative strategy, such as incorporating an additional attention mechanism, might help mitigate this issue.
2.7 Effect of Chunked Contexts
To investigate whether the models can process the full 4.096-token context cohesively, we conducted an experiment in which the serialized EHR input was divided into chunks of 512, 1.024, and 2.048 tokens. For each chunk, separate embeddings were generated and then averaged to create a final representation (see Table 5). For GTE-Qwen2-7B, the performance decrease with smaller chunks was relatively modest, indicating that the information contained within the full 4.096-token input remains effectively used even when segmented. Notably, using 512-token chunks yielded an overall AUROC performance of 0.735, compared to 0.672 when processing a contiguous 512-token context, which suggests that chunking can mitigate input constraints. In contrast, LLM2Vec-Llama-3.1-8B demonstrated improved performance with chunked inputs, consistent with its behavior on shorter context sizes. This enhancement was primarily driven by better lab value prediction, implying that LLM2Vec-Llama-3.1-8B is particularly effective when processing inputs of up to 2.048 tokens.
| GTE-Qwen2-7B | |||||
| 2 x 2.048 tokens chunks | |||||
| 4 x 1.024 tokens chunks | |||||
| 8 x 512 tokens chunks | |||||
| LLM2Vec-Llama-3.1-8B | |||||
| 2 x 2.048 tokens chunks | |||||
| 4 x 1.024 tokens chunks | |||||
| 8 x 512 tokens chunks |
| GTE-Qwen2-7B | |||||

### v3

@@ Abstract
Electronic Health Records (EHRs) offer considerable potential for clinical prediction, but their complexity and heterogeneity present significant challenges for traditional machine learning methods. Recently, domain-specific EHR foundation models trained on large volumes of unlabeled EHR data have shown improved predictive accuracy and generalization. However, their development is constrained by limited access to diverse, high-quality datasets, and by inconsistencies in coding standards and clinical practices. In this study, we explore the use of general-purpose Large Language Models (LLMs) to encode EHR into high-dimensional representations for downstream clinical prediction tasks. We convert structured EHR data into markdown-formatted plain text documents by replacing medical codes with natural language descriptions. This enables the use of LLMs and their extensive semantic understanding and generalization capabilities as effective encoders of EHRs without requiring access to private medical training data. We show that LLM-based embeddings can often match or even surpass the performance of a specialized EHR foundation model, CLMBR-T-Base, across 15 diverse clinical tasks from the EHRSHOT benchmark. To demonstrate generalizability, we further evaluate the approach on the UK Biobank (UKB) cohort, a population distinct from that used to train CLMBR-T-Base. Notably, one of the tested LLM-based models achieves superior performance for disease onset, hospitalization, and mortality prediction, highlighting robustness to shifts in patient populations. Our findings suggest that repurposed general-purpose LLMs for EHR encoding provide a scalable and generalizable alternative to domain-specific models for clinical prediction.
keywords:
electronic health records, clinical prediction, machine learning, large language models, foundation models1 Introduction
@@
EHRs are now widely used in modern healthcare, providing comprehensive, longitudinal views of a patient’s health status [1]. Machine learning methods can use this rich data for risk stratification and to support clinical decision-making [2, 3, 4]. In recent years, researchers have explored a variety of prediction tasks based on EHRs, including hospital readmission [5, 6], length of hospital stay [6], sepsis onset detection [7, 8], mortality prediction [6, 9], discharge diagnoses [6], and heart failure outcomes [10]. The overarching goal is to harness EHR data using machine learning to improve clinical outcomes and reduce healthcare costs.
@@
However, machine learning on EHR data poses significant challenges due to its inherent complexity. EHR data is characterized by variable-length sequences of patient visits, irregular sampling intervals, missing entries, heterogeneous and noisy information, and a wide range of hierarchical medical concepts [11]. As a result, deep learning models often achieve only modest improvements over traditional methods such as logistic regression or tree-based methods [6, 12, 13]. To mitigate these issues, recent approaches have employed large-scale foundation models that are pre-trained on unlabeled EHR data using unsupervised learning [14]. Many of these models adopt strategies from natural language processing, such as masked-word prediction as in BERT [15] or autoregressive next-word prediction as in GPT [16]. Treating EHR data as sequences of medical codes enables analogous methods such as masked code prediction [17, 18, 12, 19] or next code prediction [13]. However, these techniques also face significant limitations: coding standards and healthcare practices differ strongly across sites, and interoperable EHR foundation models would likely need to be trained on a wide variety of EHR datasets, which is difficult to achieve due to the sensitivity of healthcare data. Therefore, the development of EHR-specific foundation models remains constrained by the limited size and restricted availability of EHR data.
@@
In contrast, LLMs benefit from training on vast general-purpose text corpora and a broad range of natural language tasks [20, 21]. This extensive pre-training enables their language comprehension and allows them to capture domain-agnostic patterns that can be adapted for healthcare applications. Consequently, LLMs have demonstrated strong performance in extracting medical concepts [22], summarizing medical texts [23], and predicting medical outcomes [24] even in low-resource settings. However, most modern LLMs, such as GPT [25] or Llama [26], use a decoder-only transformer architecture, which complicates the generation of robust text representations. To overcome this limitation, recent work has introduced methods to convert decoder-only LLMs into effective embedding models for downstream prediction tasks [27, 28, 29, 30]. Additionally, these state-of-the-art models offer an increased context window, making them well-suited for handling long inputs such as serialized EHR data.
@@
In this study, we evaluate whether general-purpose LLM embedding models can effectively encode EHR data for clinical prediction [31] (see Fig. 1). To this end, we convert structured EHR records into plain text summaries that highlight key patient information at prediction time. Using two state-of-the-art LLM embedding models, GTE-Qwen2-7B-Instruct (GTE-Qwen2-7B) [32, 30] and LLM2Vec-Llama-3.1-8B-Instruct (LLM2Vec-Llama-3.1-8B) [27, 33], we generate high-dimensional EHR embeddings that serve as inputs to logistic regression classifiers across 15 clinical tasks from the EHRSHOT benchmark. We focus on the performance in few-shot settings to evaluate the generalization ability of this approach and conduct extensive ablation studies to identify the specific components that drive the LLM ’s effectiveness. Finally, we validate our approach on the UKB for predicting mortality, hospitalization and onset of 23 diseases [34] to assess its robustness and compare the generalizability with EHR-specific foundation models.
@@ 2 Results
2.1 Experimental Setup
Our primary analysis was conducted using the EHRSHOT database, which contains EHRs from 6,739 adult patients treated at Stanford Health Care and Lucile Packard Children’s Hospital between 1990 and 2023. This dataset includes 921,499 visits and over 41.6 million clinical events [31], and serves as a standardized benchmark with 15 clinical prediction tasks organized into 4 task groups, along with predefined splits and publicly available code. Table 1 summarizes cohort statistics, and task details are shown in Table 2. Further information on task definitions and preprocessing is provided in Section 4.1.
| Attribute | EHRSHOT | UK Biobank |
| Num Patients | 6,739 | 387,464 |
| Num Visits | 921,499 | 19,484,777 |
| Num Events | 41,661,637 | 72,265,684 |
| Num Female | 3,441 | 214,565 |
| Num Male | 3,298 | 172,899 |
| Age, mean SD | 59.3 17.9 | 56.78 8.11 |
| American Indian | 25 | 0 |
| Asian | 1,043 | 8,659 |
| Black | 298 | 5,751 |
| Pacific Islander | 74 | 0 |
| Unknown | 1,563 | 7,202 |
| White | 3,736 | 365,888 |
| Hispanic | 1,038 | - |
| Non-Hispanic | 5,701 | - |
@@ 2.1 Experimental Setup
For external validation, we used the UKB, a population-based cohort comprising 502,489 UK participants [35, 36]. This allowed us to evaluate the generalizability of our approach across healthcare systems, especially since the CLMBR-T-Base model was trained on data from the same hospital system as EHRSHOT. Due to structural differences between the datasets, not all EHRSHOT tasks were transferable to the UKB (see Section 4.8). We focused on predicting one-year risk of hospitalization, mortality, and the onset of 23 chronic diseases [34]. The processed UKB subset used in our study consisted of 387,464 patients, approximately 19.5 million visits, and over 72 million clinical events. Table 6 provides task descriptions and label distributions.
To prepare EHR data for use with LLMs, we serialized structured patient records into plain text formatted in Markdown, with a maximum length of 4,096 tokens, prioritizing more recent data. For EHRSHOT, serialization began with a reference prediction date to which all event dates were normalized, followed by demographics, recent values for 24 key LOINC concepts across Body Metrics, Vital Signs, and Lab Results, including units and classification (low/normal/high). These were followed by a summary of patient visits and reverse-chronological listings of visit-level events such as conditions, medications, and procedures. An example EHR text serialization is shown in Fig. 2, and a detailed explanation appears in Section 4.2. For the UKB, a similar serialization approach was used, with event dates normalized to each patient’s recruitment date and exclusion of detailed measurements due to data limitations.
We evaluated two instruction-tuned LLM embedding models, GTE-Qwen2-7B [32, 30] and LLM2Vec-Llama-3.1-8B [33, 27], using task-specific prompts (e.g., “Given a patient’s electronic healthcare record (EHR) in Markdown format, retrieve relevant passages that answer the query: has the patient anemia”, see Table 8). Our primary objective was to assess how general-purpose LLM-based embedding models, trained on public text data, perform in encoding EHRs for clinical prediction tasks. For comparison, we included CLMBR-T-Base, a 141-million-parameter autoregressive foundation model trained on 2.57 million de-identified EHRs from Stanford Medicine [13, 31]. For each model, we computed patient-level embeddings and trained a logistic regression classifier on the training split, with hyperparameters selected using a validation set. As a baseline, we also included a Gradient Boosted Machine (GBM) classifier trained on count-based features, which has previously demonstrated strong performance on EHR tasks [31]. See Section 4.4 for additional details on model architectures.
@@
| Model | Mortality prediction | Operational Outcomes (Hospitalization) | Assignment of New Diagnoses | Macro Avg. Across Task Groups |
| LLM-Embedding Models | ||||
| GTE-Qwen2-7B | ||||
| LLM2Vec-Llama-3.1-8B | ||||
| Baselines [31] | ||||
| CLMBR-T-Base | ||||
| Counts-based + GBM | ||||
2.2 General-Purpose LLM Embeddings Rival Domain-Specific EHR Models
@@
Using all available training and validation examples, the GTE-Qwen2-7B model outperformed CLMBR-T-Base in two of the four task categories: lab test result prediction and assignment of new diagnoses (see Table 3). In contrast, CLMBR-T-Base achieved higher performance on tasks related to operational outcomes and chest X-ray findings. LLM2Vec-Llama-3.1-8B embeddings showed slightly lower performance than GTE-Qwen2-7B, exceeding CLMBR-T-Base only in the assignment of new diagnoses task group. Both LLM-based models were often on par with or outperformed the GBM-based baseline, which is considered a strong baseline for EHR prediction tasks [12, 13]. To assess whether the models encode orthogonal information, we concatenated CLMBR-T-Base embeddings with LLM-based embeddings and observed substantial performance improvements for both LLM models. The combined embeddings achieved an average area under the receiver operating characteristic curve (AUROC) of 0.786 (0.762–0.811) for GTE-Qwen2-7B, up from 0.755 (0.730–0.780), and 0.773 (0.747–0.799) for LLM2Vec-Llama-3.1-8B, up from 0.734 (0.707–0.762). The smaller GTE-Qwen2-1.5B model achieved an average performance of 0.748 (0.722–0.774), only marginally lower than its larger counterpart, suggesting it is a viable lightweight alternative for EHR embeddings. In contrast, LLM2Vec-Llama-2-1.3B performed notably worse than the larger version, potentially due to limitations of the older Llama 2 architecture. Figure 3 shows a general trend that larger models tend to yield better performance across tasks. Among the evaluated models, CLMBR-T-Base remained the most parameter-efficient relative to predictive performance. Nonetheless, it is important to note that the LLM-based embedding models were not specifically optimized for structured EHR data, which may explain their performance differences.
On the UKB dataset, GTE-Qwen2-7B consistently outperformed CLMBR-T-Base across all three prediction task groups (see Table 4). LLM2Vec-Llama-3.1-8B surpassed CLMBR-T-Base only in the task group related to the assignment of new diagnoses. The performance drop of CLMBR-T-Base was anticipated, as the UKB represents an out-of-domain dataset with numerous disease codes not included in the model’s original vocabulary. Although we applied a careful mapping of UKB Observational Medical Outcomes Partnership (OMOP) concepts to those supported by CLMBR-T-Base to improve compatibility and predictive performance, many concepts remained unmapped due to CLMBR-T-Base’s limited vocabulary. This limitation underscores a key challenge in adapting domain-specific models to external datasets, even with considerable manual effort. In contrast, the LLM-based models required only a mapping from clinical concepts to their natural language descriptions, which were directly incorporated into the Markdown-based serialization. This simplified integration process substantially reduced implementation overhead and emphasized the strong generalization ability of LLM-based encodings across diverse coding systems and healthcare domains.
2.3 LLM Embeddings Achieve Strong Performance in Low-Data Regimes
@@
To evaluate model performance with limited training data, we conducted experiments in a few-shot setting using small numbers of training examples. The experiments followed the EHRSHOT task definitions [31]. Fig. 5 illustrates the aggregated AUROC across all subtasks within the four task categories for varying numbers of training examples. LLM embedding models performed strongly in the few-shot setting, suggesting effective transfer from general text pretraining to serialized EHR data. Notably, GTE-Qwen2-7B consistently outperformed CLMBR-T-Base across all training sizes for lab test result prediction and assignment of new diagnoses. The strong results for lab test predictions indicate that the inclusion of the last three lab values in the EHR serialization provides highly predictive features, which GTE-Qwen2-7B effectively encodes. In contrast, LLM2Vec-Llama-3.1-8B showed comparatively weaker performance in the few-shot regime. It only surpassed CLMBR-T-Base in a limited subset of tasks, primarily those involving the assignment of new diagnoses. Nonetheless, both LLM-based models consistently outperformed the count-based baseline across most configurations, underscoring the benefits of extensive pretraining. Consistent with previous findings [31], the largest performance improvements over the baseline were observed at intermediate training set sizes. As the number of labeled data samples increased, the relative advantage of pre-trained LLM-based models diminished, highlighting the reduced marginal gains of foundation models in high-data regimes. The area under the precision-recall curve (AUPRC) results are shown in Fig. 7, with task-specific AUROC and AUPRC results provided in Fig. 8 and Fig. 9 in the Appendix.
In contrast to EHRSHOT, the LLM-based methods outperformed CLMBR-T-Base in most prediction tasks on the UKB dataset. For the assignment of new diagnoses, which includes 23 tasks, both GTE-Qwen2-7B and LLM2Vec-Llama-3.1-8B outperformed CLMBR-T-Base across all sample sizes (see Fig. 10 in the Appendix). For hospitalization and mortality prediction, CLMBR-T-Base showed stronger performance in the very low-data regime. However, with more samples the performance of the LLM embedding models increased and GTE-Qwen2-7B outperformed CLMBR-T-Base. Figure 5 presents the mean AUROC across all task groups for both EHRSHOT and UKB, highlighting the observed differences. On EHRSHOT, GTE-Qwen2-7B generally matched CLMBR-T-Base in performance, while LLM2Vec-Llama-3.1-8B fell between these models and the count-based baseline. On the UKB dataset, LLM-based models and CLMBR-T-Base performed comparably with very limited training data of up to four examples. With additional training data, GTE-Qwen2-7B surpassed both CLMBR-T-Base and LLM2Vec-Llama-3.1-8B, while the latter closely approached the performance of CLMBR-T-Base. These findings underscore the capability of general-purpose LLM embeddings to rival or exceed the performance of domain-specific models, particularly in settings characterized by domain shift or limited labeled data. Detailed task-specific results are available in Fig. 12 and Fig. 13 in the Appendix.
@@
| generic instructions | |||||